# Distribution of lengths of closed saddle connections on moduli space of large genus translation surface

### v2

@@ 1. Introduction
Corollary 1.4.
Let ${\mathcal{H}}_{g}(2^{O(g)},1^{2g-2-2O(g)})$ be a stratum with $O(g)$ double zeros, then the result in Theorem 1.2 is also true.
@@ 4. Surgery
This section we will introduce some surgeries that, when combined, can collapse a closed saddle connection.
@@ 4.1. Open up a higher-order zero
By the surgery of collapsing this map is one-to-one and denote the opening surgery by $F_{1}$.
@@ 4.2. Move zero along closed curve and pinch
We have constructed the mapping $F_{1}$, which takes a closed saddle connection $\gamma$ at a double zero to two saddle connections $\gamma_{1}$ and $\gamma_{2}$ sharing same endpoints $p$ and $q$ which are both simple zeros. Moreover, the angles between the two saddle connections at the two simple zeros (on both sides) are all $2\pi$, and if we denote by $\ell_{\omega}(\gamma)=L$, we have $\ell_{\omega^{\prime}}(\gamma_{1})=L$, $\ell_{\omega^{\prime}}(\gamma_{2})=2L$, where $(X^{\prime},\omega^{\prime})=F_{1}(X,\omega)$. If we want to collapse $\gamma_{1}$ and $\gamma_{2}$ simultaneously, we have to make a surgery to move $q$ along $\gamma_{2}$ to adjust the length of $\gamma_{2}$.
@@
First we need to choose a special basis of $H^{1}(X,\Sigma;\mathbb{C})$, where $\Sigma=\{p,q,p_{2},\cdots,p_{g-1}\}$ is the zero set of $(X^{\prime},\omega^{\prime})=F_{1}(X,\omega)$. From Claim 1, the homology class $\gamma]=[\gamma_{1}+\gamma_{2}]$ is non-separable. So we can choose a basis of $H^{1}(X,\mathbb{C})$
$(\alpha_{1},\beta_{1},\cdots,\alpha_{g},\beta_{g}),$
where $\alpha_{1}=[\gamma]$.
@@
Let $\omega^{\prime}(\alpha_{1})=(x_{1},y_{1})$, $\omega^{\prime}(\beta_{1})=(x_{2},y_{2})$, $\omega^{\prime}(\alpha_{2})=(x_{3},y_{3})$ $\omega^{\prime}(\beta_{2})=(x_{4},y_{4})$. Now choose a relative cohomology class $\upsilon\in H^{1}(X,\Sigma;\mathbb{C})$ , which can be considered as a tangent vector in $T_{{\mathcal{H}}_{g}(1,1,\cdots,2)}(X^{\prime},\omega^{\prime})$ such that
(4.1) $\upsilon(\alpha_{1})=(-x_{1},-y_{1}),\upsilon(\alpha_{2})=(-\frac{y_{2}x_{1}}{y_{4}},\frac{x_{2}y_{1}}{x_{4}})$
and on the other basis we assign zero to $\upsilon$.
@@
Consider the curve in ${\mathcal{H}}_{g}(1,1,2,,\cdots,2)$: $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. If $\gamma_{1}$ and $\gamma_{2}$ don’t degenerate along the curve, then we obtain
$F_{2}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\to{\mathcal{H}}_{g}(1^{2},2^{g-1})$
such that $F_{2}[(X,\omega)]=F_{1}[(X,\omega)]+\upsilon$. From the construction above the resulting surface has two saddle connections in the relative homology class $\gamma_{1}$ and $\gamma_{2}$ with equal length.
4.2.2. $F_{2}$ is well-defined almost everywhere
We have defined a mapping if the saddle connections $\gamma_{1}$ and $\gamma_{2}$ are preserved along the curve $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$ on $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$. Next we will see the mapping can be defined for almost every translation surface in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
From above we know the moving surgery can be realized except for some of $\gamma_{1}$ or $\gamma_{2}$ degenerate along $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Suppose $T$ is the first time when the geodesic in $\gamma_{1}]$ can be represented by $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$, since the saddle connection is smooth on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$, the corner of $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$ must be $\pi$. But the holonomy of $pp_{i}$ is not changed on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$ This implies there exists some $pp_{i}$ such that its holonomy on $(X^{\prime},\omega^{\prime})$ has the same direction with $p_{0}p_{1}$. And by the relation of holonomy under $F_{1}$, this implies $p_{0}p_{i}$ has the same direction with the closed curve on $(X,\omega)$. Under period mapping, on each local chart this is a measure-zero set. So all such $(X,\omega)$ is a subset of measure zero in ${\mathcal{H}}_{g}(2^{g-1})$, also in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$.
@@ 4.2.3. Pinching
On the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ where $F_{2}$ is well-defined, we can cut along $\gamma_{1}+\gamma_{2}$ and denote the two copies of $p$ and $q$ by $^{+}, qq^{-}$, $^{+}, pp^{-}$, the two copies of $\gamma_{1}$ and $\gamma_{2}$ by $\gamma^{+}_{1}$, $\gamma^{-}_{1}$, $\gamma^{+}_{2}$, $\gamma^{-}_{2}$. Notice that the total angles at $^{+}, qq^{-}$, $^{+}, pp^{-}$ are all $2\pi$, which means $\gamma^{\pm}_{1}$ have same direction with $\gamma^{\pm}_{2}$, then glue $\gamma^{+}_{1}$ with $\gamma^{+}_{2}$, $\gamma^{-}_{1}$ with $\gamma^{-}_{2}$, which make $^{+}, qq^{-}$, $^{+}, pp^{-}$ become regular points. Since $\gamma_{1}+\gamma_{2}$ is non-separable, this surgery reduces the genus by one, and obtains a new abelian differential on which we mark two points $p^{+}$ and $p^{-}$ so as to find its inverse. Considering the choice of which zero is selected after applying the inverse map, we finally obtain:
$F_{3}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\to M\times{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]},$
where $M$ is the combinatorial data of choosing one from g1g-1 zeros.
@@
4.2.4. $F_{3}$ is measure-preserving.
This surgery can be inverse except for some directions which has saddle connections, which is a measure zero subset $A_{0}(X,\omega)$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$ for every surface in ${\mathcal{H}}_{g-1}(2^{g-2},0,0)$: choose a vector $\kappa$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$, consider the two rays from $p^{+}$ and $p^{-}$ whose holonomy is $\kappa$, if the rays exist, then cut the surface along the two rays to obtain two closed loci and glue them. Let
$\tilde{{\mathcal{H}}}=\{(X,\omega,\tau):(X,\omega)\in{\mathcal{H}}_{g-1}(2^{g-2},0,0),\tau\in A_{0}(X,\omega)\},$
which is a measure zero subset of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$. So we have
$F_{3}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1}))\subset[{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})]\setminus\tilde{{\mathcal{H}}}$.
Moreover, under local chart determined by the basis we choose,
$F_{3}(z_{1},\cdots,z_{n})=(z_{1},z_{i}+\upsilon(z_{i})),$
whose Jacobian has deteminant one and thus is measure-preserving.
@@ 5. Collapse closed saddle connections simultaneously
From the construction above we can collapse one closed saddle connection on some $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$ and obtain a new surface $(X^{\prime},\omega^{\prime})\in{\mathcal{H}}_{g-1}(2^{g-2},0,0)$. if we want to collapse any $k$ closed saddle connections, we have to ensure the locus of surgery for these saddle connections are disjoint. For a given interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, define the exception set $\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ to be the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ such that on every $(X,\omega)\in\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ there are two closed saddle connections with length in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ whose surgery loci intersect. First we will show that the measure of exception set goes to zero as genus goes to infinity.

### v3

@@ 1. Introduction
Moreover, we consider general stratum ${\mathcal{H}}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})$, where $\lim\limits_{g\to\infty}\frac{O_{i}(g)}{g}=c_{i}$, and we have
Theorem 1.4.
Let ${\mathcal{H}}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})$ be a stratum where $\lim\limits_{g\to\infty}\frac{O_{i}(g)}{g}=c_{i}$. Then for a interval $[a,b]$, $L_{g,[a,b]}(X,\omega)$ converges in distribution to a Poisson distributions of means $\lambda_{[a,b]}$, where
$\lambda_{[a,b]}=\sum_{i=1}^{k-1}c_{i}(m_{i}^{2}-1)\pi(b^{2}-a^{2}).$
@@ 4. Surgery
This section we will introduce some surgeries that, when combined, can collapse a closed saddle connection. Then we will use a series of such surgeries to construct a mapping to collapse a closed saddle connection.
@@ 4.2. Move zero along closed curve and pinch
We have constructed the opening surgery, which takes a closed saddle connection $\gamma$ at a double zero to two saddle connections $\gamma_{1}$ and $\gamma_{2}$ sharing same endpoints $p$ and $q$ which are both simple zeros. Moreover, the angles between the two saddle connections at the two simple zeros (on both sides) are all $2\pi$, and if we denote by $\ell_{\omega}(\gamma)=L$, we have $\ell_{\omega^{\prime}}(\gamma_{1})=L$, $\ell_{\omega^{\prime}}(\gamma_{2})=2L$, where $(X^{\prime},\omega^{\prime})=F_{1}(X,\omega)$. If we want to collapse $\gamma_{1}$ and $\gamma_{2}$ simultaneously, we have to make a surgery to move $q$ along $\gamma_{2}$ to adjust the length of $\gamma_{2}$.
@@
Let $(X^{\prime},\omega^{\prime})$ be the resulting surface of $(X,\omega)$ after opening surgery for some closed saddle connection $\gamma$. First we need to choose a special basis of $H^{1}(X,\Sigma;\mathbb{C})$, where $\Sigma=\{p,q,p_{2},\cdots,p_{g-1}\}$ is the zero set of $(X^{\prime},\omega^{\prime})$. From Claim 1, the homology class $\gamma]=[\gamma_{1}+\gamma_{2}]$ is non-separable. Let $\alpha_{1}=[\gamma]$, $\beta_{1}$ be the homology class dual to $\alpha_{1}$. Then choose $\beta_{2}$ to be the homology class of curves whose flat geodesic has minimal vertical projection during all the simple, non-separable closed curves on $X\setminus\alpha_{1}$. Note that the set of those flat surfaces which have horizontal geodesic or two geodesics with same vertical length has measure zero, we can suppose the minimal is not zero. Let $\alpha_{2}$ be the homology class dual to $\beta_{2}$. Now we can extend the curves we have chosen to be a canonical basis of $H^{1}(X,\mathbb{C})$
$(\alpha_{1},\beta_{1},\cdots,\alpha_{g},\beta_{g}).$
@@
Let $\omega^{\prime}(\alpha_{1})=(x_{1},y_{1})$, $\omega^{\prime}(\beta_{1})=(x_{2},y_{2})$, $\omega^{\prime}(\alpha_{2})=(x_{3},y_{3})$ $\omega^{\prime}(\beta_{2})=(x_{4},y_{4})$ and suppose $x_{1}y_{2}-x_{2}y_{1}>0,y_{4}>0$. Now choose a relative cohomology class $\upsilon\in H^{1}(X,\Sigma;\mathbb{C})$ , which can be considered as a tangent vector in $T_{{\mathcal{H}}_{g}(1,1,\cdots,2)}(X^{\prime},\omega^{\prime})$ such that
(4.1) $\upsilon(\alpha_{1})=(-x_{1},-y_{1}),\upsilon(\alpha_{2})=(\frac{y_{2}x_{1}}{y_{4}}-\frac{x_{2}y_{1}}{y_{4}},0)$
and on the other basis we assign zero to $\upsilon$. Locally from period mapping $\upsilon$ is a smooth vector field.
@@
Consider the curve in ${\mathcal{H}}_{g}(1,1,2,,\cdots,2)$: $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. If $\gamma_{1}$ and $\gamma_{2}$ don’t degenerate along the curve, then we obtain a new surface $(X^{\prime},\omega^{\prime})+\upsilon$. Since the vertical projection of all saddle connections is not changed, so do all of simple non-separable closed curves. Then $\beta_{2}$ is also the simple non-separable closed curve whose vertical projection is minimal along $(X^{\prime},\omega^{\prime})+t\upsilon$. So when a closed saddle connection $\gamma$ is chosen, at any time $t_{0}$ we can always choose $\beta_{2}$ and $\upsilon$, which guarantees $\upsilon$ is indepedent to the locally period mapping and make it a vector field globally. From the construction above the resulting surface has two saddle connections in the relative homology class $\gamma_{1}$ and $\gamma_{2}$ with equal length.
4.2.2. moving surgery is well-defined almost everywhere
Let $(X,\omega)$ be a surface in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. After opening a closed saddle connection $\gamma$, $(X,\omega)$ becomes $(X^{\prime},\omega^{\prime})$. We have defined the moving surgery if the saddle connections $\gamma_{1}$ and $\gamma_{2}$ are preserved along the curve $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Next we will see this surgery can be defined for almost every translation surface in the set of resulting surfaces under opening surgery on $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
From above we know the moving surgery can be realized except for some of $\gamma_{1}$ or $\gamma_{2}$ degenerate along $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Suppose $T$ is the first time when the geodesic in $\gamma_{1}]$ can be represented by $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$, since the saddle connection is smooth on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$, we must have
$\vec{pp_{i}}+\vec{p_{i}p_{j}}+\cdots+\vec{p_{k}q}=\omega^{\prime}+T\upsilon(\gamma_{1}),$
$|pp_{i}|+\cdots+|p_{k}q|=|\gamma_{1}|.$
So the corner of $pp_{i}-p_{i}p_{j}-\cdots-\vec{p_{k}q}$ must be $\pi$. But the holonomy of $pp_{i}$ is not changed on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$ This implies there exists some $pp_{i}$ such that its holonomy on $(X^{\prime},\omega^{\prime})$ has the same direction with $p_{0}p_{1}$. And by the relation of holonomy under opening surgery, this implies $p_{0}p_{i}$ has the same direction with the closed curve on $(X,\omega)$. Under period mapping, on each local chart this is a measure-zero set. So all such $(X,\omega)$ is a subset of measure zero in ${\mathcal{H}}_{g}(2^{g-1})$, also in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$.
Moreover, if some $(Y,\theta)$ is a resulting surface of moving surgery, note that there exists at least one simple closed curve consisting of a pair of non-homologous saddle connections sharing same endpoints, which we denote by $\alpha_{1}$. By the construction of moving surgery, the vertical length of $\beta_{2}$ is fixed and minimal along $(X,\omega)+t\upsilon$. Then we can choose the simple, non-separable closed curve whose vertical length is minimal on $Y\setminus\alpha_{1}$ and denote it by $\beta_{2}$, then construct the tangent vector $\upsilon$, where $\alpha_{1}$ is a homology class consists of two non-homologous saddle connections which has same endpoints and same lengths, since $(Y,\theta)$ is in the image, there exists at least one pair of such saddle connections. And from the construction if $(Y,\theta)$ has two pre-image it must have at least two such pairs of saddle connections.
@@ 4.2.3. Pinching
On the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ where the moving surgery is well-defined, we can cut along $\gamma_{1}+\gamma_{2}$ and denote the two copies of $p$ and $q$ by $^{+}, qq^{-}$, $^{+}, pp^{-}$, the two copies of $\gamma_{1}$ and $\gamma_{2}$ by $\gamma^{+}_{1}$, $\gamma^{-}_{1}$, $\gamma^{+}_{2}$, $\gamma^{-}_{2}$. Notice that the total angles at $^{+}, qq^{-}$, $^{+}, pp^{-}$ are all $2\pi$, which means $\gamma^{\pm}_{1}$ have same direction with $\gamma^{\pm}_{2}$, then glue $\gamma^{+}_{1}$ with $\gamma^{+}_{2}$, $\gamma^{-}_{1}$ with $\gamma^{-}_{2}$, which make $^{+}, qq^{-}$, $^{+}, pp^{-}$ become regular points. Since $\gamma_{1}+\gamma_{2}$ is non-separable, this surgery reduces the genus by one, and obtains a new abelian differential on which we mark two points $p^{+}$ and $p^{-}$ so as to find its inverse.
@@
4.2.4. collapsing a closed saddle connection
Now denote by
$\tilde{\mathcal{C}}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$
to be the set of $(X,\omega,\gamma)$, where $(X,\omega)\in\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ and $\gamma$ is a closed saddle connection on $(X,\omega)$ with length in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Then for $(X,\omega,\gamma)\in\tilde{\mathcal{C}}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, we make opening, moving and pinching surgeries in sequence and get a new surface in ${\mathcal{H}}_{g-1}(2^{g-1},0,0)$ and a vector in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$. This defines a mapping
$F:\tilde{\mathcal{C}}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\to{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}.$
Next we will show the mapping is one to one and measure-preserving almost everywhere.
Proposition 4.4.
$F$ is invertible except for a subset of measure zero of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$.
Proof.
Let $(X,\omega),\vec{a}]\in{\mathcal{H}}_{g-1}(2^{g-1},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$. Since the opening surgery is inversed, it suffices to consider when the pinching and moving sugeries can be inversed.
For pinching surgery, it can be inversed except for some directions which has saddle connections, which is a measure zero subset $A_{0}(X,\omega)$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$ for every surface in ${\mathcal{H}}_{g-1}(2^{g-2},0,0)$: choose a vector $\kappa$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$, consider the two rays from $p^{+}$ and $p^{-}$ whose holonomy is $\kappa$, if the rays exist, then cut the surface along the two rays to obtain two closed loci and glue them.
Suppose after the inverse of pinching $(X,\omega),\vec{a}]$ becomes $(X^{\prime},\omega^{\prime},\gamma)$, where $\gamma$ is a simple closed curve consisting of two saddle connections with same length. For moving surgery, from 4.2.2, it can be inversed except for $(X^{\prime},\omega^{\prime})$ having the other closed curve $\gamma^{\prime}$ consisting of a pair of non-homologous saddle connections which have same length. But $\gamma^{\prime}$ is fixed during the pinching surgery and other saddle connections are not changed, this means $(X,\omega)$ having a simple closed curve consisting of two saddle connections with same length, which is a measure zero set of ${\mathcal{H}}_{g-1}(2^{g-1},0,0)$.
So for any $(X,\omega),\vec{a}]\in{\mathcal{H}}_{g-1}(2^{g-1},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$ except for a subset of measure zero, the surgeries can be inversed, which means $F$ is invertible almost everywhere. ∎
Now we will see the mapping is also measure-preserving on . In fact, under local chart determined by the basis we choose,
$F(z_{1},\cdots,z_{n})=(z_{1},z_{i}+\upsilon(z_{i})),$
whose Jacobian has deteminant one. Since Masur-Veech measure is invariant under coordinate transformation of period mapping, $F$ is measure-preserving locally. Let $\tilde{{\mathcal{H}}}$ be the subset on which $F$ is invertible which is a full measure subset of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$. Then we have
$\tilde{{\mathcal{H}}}\subset F(\mathcal{C}^{1,0}_{g,Per}(2^{g-1}))\subset{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}.$
[…]
$)\operatorname{Vol}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}])=\operatorname{Vol}(F^{-1}(\tilde{{\mathcal{H}}}$))
[…]
$\operatorname{Vol}(F^{-1}(\tilde{{\mathcal{H}}}))=\operatorname{Vol}(\tilde{{\mathcal{H}}})=\operatorname{Vol}([{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})])$
. Combine the two equality above,
$\operatorname{Vol}([{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})])=\operatorname{Vol}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]).$
@@ 5. Collapse closed saddle connections simultaneously
This section we will prove Theorem 1.2.
From the construction above we can collapse one closed saddle connection on $(X,\omega)$ in a full measure set of $\mathcal{C}^{1,0}_{g}(2^{g-1})$ and obtain a new surface $(X^{\prime},\omega^{\prime})\in{\mathcal{H}}_{g-1}(2^{g-2},0,0)$. if we want to collapse any $k$ closed saddle connections, we have to ensure the locus of surgery for these saddle connections are disjoint. For a given interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, define the exception set $\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ to be the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ such that on every $(X,\omega)\in\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ there are two closed saddle connections with length in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ whose surgery loci intersect. First we will show that the measure of exception set goes to zero as genus goes to infinity.